OpenStudy (yb1996):

Find the limit as h approaches 0 of (f(x+h)-f(x-h))/(2h).

OpenStudy (anonymous):

any particular \(f(x)\)?

OpenStudy (yb1996):

no

OpenStudy (anonymous):

then it could be equal to anything, or not even exist at all

OpenStudy (anonymous):

but perhaps you are supposed to look at it is \[\frac{1}{2}\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

do you recognize the expression after the \(\frac{1}{2}\) ?

OpenStudy (yb1996):

Yes, that would just be f'(x)/2

OpenStudy (anonymous):

yes

OpenStudy (yb1996):

But if you are going to do it that way, then it would be written \[\frac{ 1 }{ 2 }[\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }+\lim_{h \rightarrow 0}\frac{ f(x-h)+f(x) }{ h }]\]

OpenStudy (anonymous):

oooh i read it wrong, sorry

OpenStudy (yb1996):

no problem, I'm just a bit confused as to how to get a general solution to the problem

OpenStudy (anonymous):

what you wrote is the derivative twice

OpenStudy (yb1996):

wait...really?

OpenStudy (yb1996):

So the second limit is also a derivative?

OpenStudy (anonymous):

i think so, let me check more carefully hold on yes the second is also since it is a two sided limit \(h\) has to go to zero from both directions so there is no difference between them

OpenStudy (anonymous):

oh no wrong again, let me pay better attension the second is plus not minus

OpenStudy (anonymous):

but if the function is continuous then \[\lim_{h\to 0}f(x-h)=f(x)\]

OpenStudy (anonymous):

i better do this with pencil and paper so as not to mess you about

OpenStudy (yb1996):

ok, I'm not told that the function is continuous, and I'm not really sure whether it would be proper to infer that, but yes, it would make the problem a whole lot easier if it was continuous

OpenStudy (anonymous):

if it isn't then you can't say anything

OpenStudy (yb1996):

So, would the answer just be f'(x)/2 ?

OpenStudy (anonymous):

no just \(f'\) since you have two of them and you are dividing by 2 lets go carefully and make sure it it is clear and also correct

OpenStudy (anonymous):

you already broke it apart by adding and subtracting \(f(x)\) to get \[\frac{ 1 }{ 2 }[\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }+\lim_{h \rightarrow 0}\frac{ f(x-h)+f(x) }{ h }]\]

OpenStudy (anonymous):

oops typo there, it should be \[\frac{ 1 }{ 2 }[\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h }+\lim_{h \rightarrow 0}\frac{ -f(x-h)+f(x) }{ h }]\]

OpenStudy (anonymous):

the first part is evidently the derivative right?

OpenStudy (yb1996):

yeah

OpenStudy (anonymous):

to see that the second one is also replace \(h\) by \(-h\)

OpenStudy (anonymous):

let me know if that is clear or not

OpenStudy (yb1996):

Yeah, that makes sense, because -h would turn the second limit into the second limit. But if you replace h by -h for the second limit, wouldn't you have to do that with the first one as well?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

taking the limits separately

OpenStudy (yb1996):

ok, and I have another question: would replacing h with -h mean that h would have to be negative for the limit to work?

OpenStudy (anonymous):

no h goes to zero from both directions it is just an algebra trick to turn \[\frac{-f(x-h)+f(x)}{h}\] in to \[\frac{f(x+h)-f(x)}{h}\]

OpenStudy (yb1996):

ok, that makes sense. Thank you for your help!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

sorry i screwed up so much before reading properly

OpenStudy (yb1996):

don't worry, we all make mistakes